OpenStudy (anonymous):

If 8 percent of c is equal to 12 percent of d, which of the following is equal to c+d?

OpenStudy (anonymous):

.08 c = .12d

OpenStudy (anonymous):

8c/100=12d/100

OpenStudy (anonymous):

so of means mulitplication?

OpenStudy (anonymous):

Exactly!

OpenStudy (anonymous):

2c=3d c=k/2 d=k/3 c+d=(5k)/6

OpenStudy (anonymous):

answer choices are 1.5d,2d,2,5d,3d, and 5d

OpenStudy (anonymous):

you got that from crossmultiplyng?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

where did you get k from?

OpenStudy (anonymous):

2.5d is answer

OpenStudy (anonymous):

so let me get this straight, 2c=3d

OpenStudy (anonymous):

c=3d/2

OpenStudy (anonymous):

it is just constant..

OpenStudy (anonymous):

so

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

c=1.5

OpenStudy (anonymous):

1.5d specifically

OpenStudy (anonymous):

so 1.5d plus d equals 2.5

OpenStudy (anonymous):

which is what you stated

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

since c=3d/2

OpenStudy (anonymous):

c=1.5d

OpenStudy (anonymous):

so, 1.5d plus d=2.5d

OpenStudy (anonymous):

which is the same as c+d

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

since c=1.5d

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

C +d = 3d/2 + d = 1.5 d + d

OpenStudy (anonymous):

ty